Quinine (a non-volatile, non electrolyte) has the empirical formula C10H12NO.A sample of

quinine (0.162 g) was mixed with camphor (10.0 g) and the freezing point of the mixture was found to be 175.2 oC. The freezing point of pure camphor was determined to be 177.2 oC.

[Kf (camphor) = 40.0 oC Kg mol-1]

(a) Calculate the freezing point depression of the solution.

(b) Calculate the molality of the quinine.

(c) Calculate the moles of quinine.

(d) Calculate the molar mass of quinine

(e) What is the molecular formula of quinine?


1
Expert's answer
2022-04-22T08:20:03-0400

The molecular formula of quinine is C20H24N2O2

The molar mass of quinine: 12.01 * 20 + 1.01 * 24 + 14.01 * 2 + 16.00 * 2 = 324.42 g/mol

The moles of quinine: 0.162 / 324.42 = 0.0005 mol

The molality: 0.0005 / 0.01 = 0.05 mol/kg

The freezing point depression of the solution: ΔT = Kf*m = 40.0 * 0.05 = 2.0 0C